OpenStudy (amtran_bus):

Just wanted to double check my answer, will take 2 seconds.

OpenStudy (amtran_bus):

9.71 is what percent of 9.81

OpenStudy (anonymous):

(9.71/9.81)*100

OpenStudy (amtran_bus):

Thought is was 1%.

OpenStudy (amtran_bus):

!!

OpenStudy (anonymous):

Do the is over of thing meaning make it into a fraction. by taking is over of then and equal sign and then the number over 100 but you would put number is the is over of spots.. it really helps that's how i started to understand it better (:!!!

OpenStudy (stamp):

9.71 is what percent of 9.81 9.71 is x percent of 9.81 9.71 is x % of 9.81 9.71 = x % 9.81 9.71 = ( x / 100 ) 9.81 9.71 / 9.81 = x / 100 x = ( 9.71 / 9.81 ) * 100 = http://www.wolframalpha.com/input/?i=%28+9.71+%2F+9.81+%29+*+100 = 98.9806 so: 9.71 is 98.9806 percent of 9.81.